# Diffraction by a Right-Angle Penetrable Wedge: Closed-Form Solution for $\nu=\sqrt{2}$

### v2

@@ 10.4 Canonical no-double-counting representation
Assume the analytic strip and growth framework of §1.1 and the uniqueness principle of Lemma 1.3. Let the forcing pole set $\{u_{\ell}\}$ be as in §4. Let the coefficient tables $(\alpha_{\ell})$, $(\beta_{\ell})$ and $(C_{\ell})$ be given by Propositions 6.16.3, and define $(d_{\ell})$ by Proposition 10.1.
$Q_{\mathrm{scat}}(u)=\sum_{\ell\in I_{\mathrm{scat}}}C_{\ell}\Bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\Bigr]+R(u),$ (77)
where $C_{\ell}$ is given by Proposition 6.3 and $R$ is as above. Then $Q_{\mathrm{scat}}$ has poles exactly at the forcing points $u_{\ell}$ ($\ell\in I_{\mathrm{scat}}$) with residues $C_{\ell}$, and the remainder $R$ is pole-free at every $u_{\ell}$ and analytic at $u_{0}$. In particular, the decomposition contains no double counting of principal parts from the singular channel $P_{1,3}$.
@@
Assume the analytic and growth hypotheses on Sommerfeld densities from Lemma 1.3, together with the gauge normalization $Q_{\mathrm{scat}}(u_{0})=0$. Let $Q_{\mathrm{scat}}$ be the scattered spectral density constructed in Theorem 10.4. If $\widetilde{Q}_{\mathrm{scat}}$ is any other meromorphic function on the physical branch with at most simple poles at the forcing points $\{u_{\ell}:\ \ell\in I_{\mathrm{scat}}\}$, satisfying the same spectral functional system and scattered allocation, and the same gauge normalization $\widetilde{Q}_{\mathrm{scat}}(u_{0})=0$, then $\widetilde{Q}_{\mathrm{scat}}\equiv Q_{\mathrm{scat}}$. Equivalently, within this Sommerfeld class the corresponding scattered fields are unique.
@@
Let $H:=\widetilde{Q}_{\mathrm{scat}}-Q_{\mathrm{scat}}$. By linearity of the functional system, $H$ satisfies the associated homogeneous system (zero forcing), and by the local residue relations in Appendix A its principal parts at each forcing point are uniquely determined. Since both $\widetilde{Q}_{\mathrm{scat}}$ and $Q_{\mathrm{scat}}$ satisfy the same residue tables (Propositions 6.16.3), the difference $H$ is analytic at every $u=u_{\ell}$ ($\ell\in I_{\mathrm{scat}}$). Moreover, $H$ has no jump across the contour system defining the additive Riemann–Hilbert problem, so the uniformization $u\mapsto(t,Y)$ implies that $H$ extends to a holomorphic elliptic function on the square torus. A holomorphic elliptic function is constant, hence $H\equiv c$. Finally, the gauge normalization gives $c=H(u_{0})=0$, so $H\equiv 0$. If the gauge is not imposed then $c$ is the only remaining ambiguity; this constant does not affect the Sommerfeld integrals because they involve the difference $Q(\theta+z)-Q(\theta-z)$ (cf. Lemma 1.3). ∎
@@ 11 Analyticity at the incident spectral point
Assume the analytic strip and growth framework of §1.1 and the uniqueness principle of Lemma 1.3. Let the forcing pole set $\{u_{\ell}\}$ be as in §4, and impose the scattered allocation by excluding the incident label $\ell_{\mathrm{inc}}=(0,+,-)$ from the inside set, i.e., $I_{\mathrm{scat}}=I\setminus\{\ell_{\mathrm{inc}}\}$. Then the scattered spectral density $Q_{\mathrm{scat}}(\zeta)$ is analytic at the incident spectral point $\zeta=\zeta_{i}=\theta_{i}+i\varepsilon$ (for each fixed $\varepsilon>0$).
@@
By Theorem 10.4, the function $Q_{\mathrm{scat}}$ admits the canonical decomposition
$Q_{\mathrm{scat}}(u)=\sum_{\ell\in I_{\mathrm{scat}}}C_{\ell}\bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\bigr]+R(u),$
where each zeta difference has a simple pole only at $u=u_{\ell}$ and the remainder $R$ is pole-free at every forcing point. Since $\ell_{\mathrm{inc}}\notin I_{\mathrm{scat}}$, no term in the zeta sum has a pole at the incident point $u=u_{\ell_{\mathrm{inc}}}$. Moreover, the definitions of $A(u)$, $B(u)$ and hence of $P_{1,3}(u)$ and $R(u)$ involve sums only over $I_{\mathrm{scat}}$, so $R(u)$ is analytic at $u=u_{\ell_{\mathrm{inc}}}$ as well. Therefore $Q_{\mathrm{scat}}$ is analytic at $u=u_{\ell_{\mathrm{inc}}}$, and hence, by the physical lift $u=u(\zeta)$, analytic at $\zeta=\zeta_{i}$. ∎
@@ 12 Radiation condition and Meixner edge condition in the spectral formulation
$\mathcal{S}_{\eta}:=\{\zeta\in\mathbb{C}:|\Im\zeta|<\eta\},$ (78)
@@
$|Q(\zeta)|+|S(\zeta)|\leq C(1+|\zeta|)^{N},\qquad\zeta\in\mathcal{S}_{\eta}\setminus\{\text{poles}\}.$ (79)
@@
$Q_{\mathrm{scat}}(u_{0})=0,\qquad\text{equivalently }\quad Q_{\mathrm{scat}}(\zeta)\to 0\ \text{as }\Im\zeta\to+\infty,$ (80)
(and likewise for $S_{\mathrm{scat}}$).
Then the Sommerfeld integrals (2)–(3) define classical solutions of the Helmholtz equations in their respective sectors, satisfy the Sommerfeld radiation condition as $r\to\infty$, and satisfy the Meixner finite-energy condition at the wedge apex $r\to 0$.
@@
For $r\to\infty$, steepest descent on the phase $\cos z$ along the deformed contour yields an outgoing leading term proportional to $e^{ikr}/\sqrt{r}$, with remainder $o(r^{-1/2})$; see, for example, [17, Ch. 6] or [18, § 2.4] for general stationary-phase estimates, and [3] for the diffraction interpretation in wedge geometries. The outgoing far-field expansion implies the Sommerfeld radiation condition.
For $r\to 0$, one expands $e^{ikr\cos z}=1+O(r)$ uniformly on $\gamma$ and uses (80) together with the strip analyticity to shift the contour upward, obtaining boundedness of $u$ and its first derivatives in a neighborhood of the apex; boundedness of $u$ and $\nabla u$ implies the Meixner finite-energy condition (see [6, § 3] and the discussion in [4, Ch. 3]). ∎
Corollary 12.2 (Radiation and Meixner for the constructed solution).
@@
The only singularities of $Q_{\mathrm{scat}}$ and $S_{\mathrm{scat}}$ in the strip are the prescribed simple poles (with $\varepsilon>0$ displacing them away from $\gamma$), and $Q_{\mathrm{scat}}$, $S_{\mathrm{scat}}$ are $2\pi$-periodic in $\Re\zeta$ away from poles because $s_{\zeta}=e^{i(\zeta-\theta_{w})}$ is $2\pi$-periodic. Hence the growth condition (79) holds (in fact with $N=0$) on compact subsets of the strip avoiding the poles. Proposition 12.1 applies. ∎
@@ 13 Far-field diffraction coefficient
The contour deformation and stationary-phase evaluation used in Section 13 are justified under standard analyticity and growth hypotheses for $Q_{\mathrm{scat}}(\zeta)$ in a strip containing the real axis; see, for example, Bleistein–Handelsman [17, §2.4] or Wong [18, Ch. II]. In the present lemniscatic case these hypotheses are met for $\varepsilon>0$ because $Q_{\mathrm{scat}}$ is given by an explicit elliptic-function representation (Theorem 10.4) and the forcing poles are displaced off the real $\zeta$–axis by the limiting absorption parameter.
@@
Assume the analytic strip and growth framework of §1.1 and the uniqueness principle of Lemma 1.3. Let the forcing pole set $\{u_{\ell}\}$ be as in §4. Let the coefficient tables $(\alpha_{\ell})$, $(\beta_{\ell})$ and $(C_{\ell})$ be given by Propositions 6.16.3, and define $(d_{\ell})$ by Proposition 10.1.
$\boxed{\ D(\theta,\theta_{i})=e^{-i3\pi/4}\sqrt{\frac{2}{\pi k_{0}}}\ Q_{\mathrm{scat}}(\theta)\ }.$ (81)
Here $Q_{\mathrm{scat}}(\theta)$ denotes the physical boundary value of $Q_{\mathrm{scat}}(\zeta)$ at $\zeta=\theta$, obtained by evaluating the physical lift $u(\zeta)$ and taking the limiting absorption limit $\varepsilon\to 0^{+}$ at the end.
@@
$u_{\mathrm{scat}}^{(0)}(r,\theta)=\frac{1}{2\pi i}\int_{\gamma}e^{ik_{0}r\cos z}\Bigl(Q_{\mathrm{scat}}(\theta+z)-Q_{\mathrm{scat}}(\theta-z)\Bigr)\,dz.$ (82)
@@
$u_{\mathrm{scat}}^{(0)}(r,\theta)=\frac{1}{2\pi i}\left(\int_{\gamma}+\int_{-\gamma}\right)e^{ik_{0}r\cos z}\,Q_{\mathrm{scat}}(\theta+z)\,dz.$ (83)
Under the analyticity and strip-growth hypotheses stated above, the union $\gamma\cup(-\gamma)$ may be deformed to the steepest descent rays through the saddle $z=0$, where $\cos z=1-\tfrac{1}{2}z^{2}+O(z^{4})$. The leading contribution is therefore
$u_{\mathrm{scat}}^{(0)}(r,\theta)\sim\frac{2}{2\pi i}\,Q_{\mathrm{scat}}(\theta)\,e^{ik_{0}r}\int_{-\infty}^{\infty}\exp\!\left(-i\frac{k_{0}r}{2}x^{2}\right)\,\mathrm{d}x,\qquad r\to\infty.$ (84)
Using $\int_{-\infty}^{\infty}e^{-iax^{2}}\,\mathrm{d}x=\sqrt{\pi/a}\,e^{-i\pi/4}$ for $a>0$ yields
$u_{\mathrm{scat}}^{(0)}(r,\theta)\sim e^{ik_{0}r}\,\frac{D(\theta,\theta_{i})}{\sqrt{r}},\qquad D(\theta,\theta_{i})=e^{-i3\pi/4}\sqrt{\frac{2}{\pi k_{0}}}\,Q_{\mathrm{scat}}(\theta),$ (85)
which is the claimed diffraction coefficient. ∎
This normalization is consistent with the standard two-dimensional GTD convention for wedge diffraction [3].
@@ 14 Conclusion and outlook
  • A Sommerfeld spectral representation for the transmission problem is reduced to a closed two-face functional system in the spectral variable (Section 1.1).
@@
  • The jet-killing polynomials $p,q$ are constructed so that $A(u)+p(t(u))=\mathcal{O}(t(u)^{4})$ and $B(u)+q(t(u))=\mathcal{O}(t(u)^{4})$ on $\Omega_{+}^{\mathrm{phys}}$, yielding analyticity at the physical basepoint $u_{0}$ (Section 8).
@@ 15 Symbolic evaluation recipe
    Define $u_{\ell}$ (physical lift) by the uniformization $\wp(u_{\ell})=x_{W}(t_{\ell})$ and $\tfrac{1}{2}\wp^{\prime}(u_{\ell})=y_{W}(t_{\ell},Y_{\ell})$.
@@
    Compute $\alpha_{\ell}$, $\beta_{\ell}$, and $C_{\ell}$ from Proposition 6.1, 6.2, and 6.3.
@@
    Compute jet coefficients $(p_{1},p_{2},p_{3})$ and $(q_{1},q_{2},q_{3})$ from the defining sums, or equivalently by summing the canonical per-pole summands in Propositions 6.46.5 evaluated at $(t_{\ell},Y_{\ell})$.
@@
    Build $A(u)$ and $B(u)$ from their definitions.
@@
    Build $P_{1,3}(u)$ and compute $d_{\ell}$ from (75) or Proposition 10.1. Then form $R(u)$.
@@
    For a given $\zeta$, compute $(t(\zeta),Y(\zeta))$ from the physical branch of $s(t,Y)=s_{\zeta}$, lift to $u(\zeta)$ via the uniformization, and evaluate $Q_{\mathrm{scat}}(\zeta)=Q_{\mathrm{scat}}(u(\zeta))$.
@@
    Obtain the far-field diffraction coefficient from the diffraction theorem above.

### v3

@@ 10.4 Canonical no-double-counting representation
Assume the analytic strip and growth framework of §1.1 and the uniqueness principle of Lemma 1.3. Let the forcing pole set $\{u_{\ell}\}$ be as in §4. Let the coefficient tables $(\alpha_{\ell})$, $(\beta_{\ell})$ and $(C_{\ell})$ be given by Propositions 6.16.3, and define $(d_{\ell})$ by Proposition 10.1. Set
$Q_{\mathrm{scat}}(u)=\sum_{\ell\in I_{\mathrm{scat}}}C_{\ell}\big[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\big]+R(u),$ (67)
where $R$ is defined by (66). Then $Q_{\mathrm{scat}}$ has poles exactly at the forcing points $u_{\ell}$ ($\ell\in I_{\mathrm{scat}}$) with residues $C_{\ell}$, and the remainder $R$ is pole-free at every $u_{\ell}$ and analytic at $u_{0}$. In particular, the decomposition contains no double counting of principal parts from the singular channel $P_{1,3}$.
@@
Assume the analytic and growth hypotheses on Sommerfeld densities from Lemma 1.3, together with the gauge normalization $Q_{\mathrm{scat}}(u_{0})=0$. Let $Q_{\mathrm{scat}}$ be the scattered spectral density constructed in Theorem 10.4. If $\widetilde{Q}_{\mathrm{scat}}$ is any other meromorphic function on the physical branch with at most simple poles at the forcing points $\{u_{\ell}:\ell\in I_{\mathrm{scat}}\}$, satisfying the same spectral functional system and scattered allocation, and the same gauge normalization $\widetilde{Q}_{\mathrm{scat}}(u_{0})=0$, then $\widetilde{Q}_{\mathrm{scat}}\equiv Q_{\mathrm{scat}}$.
@@
Let $H:=\widetilde{Q}_{\mathrm{scat}}-Q_{\mathrm{scat}}$. By linearity of the functional system, $H$ satisfies the associated homogeneous system (zero forcing), and by the local residue relations in Appendix A its principal parts at each forcing point are uniquely determined. Since both $\widetilde{Q}_{\mathrm{scat}}$ and $Q_{\mathrm{scat}}$ satisfy the same residue tables (Propositions 6.16.3), the difference $H$ is analytic at every $u=u_{\ell}$ ($\ell\in I_{\mathrm{scat}}$). Moreover, $H$ has no jump across the contour system defining the additive Riemann–Hilbert problem, so the uniformization $u\mapsto(t,Y)$ implies that $H$ extends to a holomorphic elliptic function on the square torus. A holomorphic elliptic function is constant, hence $H\equiv c$. Finally, the gauge normalization gives $c=H(u_{0})=0$, so $H\equiv 0$. If the gauge is not imposed then $c$ is the only remaining ambiguity; this constant does not affect the Sommerfeld integrals because they involve the difference $Q(\theta+z)-Q(\theta-z)$. ∎
@@ 11 Analyticity at the incident spectral point
Assume the analytic strip and growth framework of §1.1 and the uniqueness principle of Lemma 1.3. Let the forcing pole set $\{u_{\ell}\}$ be as in §4, and impose the scattered allocation by excluding the incident label $\ell_{\mathrm{inc}}=(0,+,-)$ from the inside set, i.e. $I_{\mathrm{scat}}=I\setminus\{\ell_{\mathrm{inc}}\}$. Then the scattered spectral density $Q_{\mathrm{scat}}(\zeta)$ is analytic at the incident spectral point $\zeta=\zeta_{i}=\theta_{i}+\mathrm{i}\varepsilon$ (for each fixed $\varepsilon>0$).
@@
By Theorem 10.4, the function $Q_{\mathrm{scat}}$ admits the canonical decomposition (67), where each zeta difference has a simple pole only at $u=u_{\ell}$ and the remainder $R$ is pole-free at every forcing point. Since $\ell_{\mathrm{inc}}\notin I_{\mathrm{scat}}$, no term in the zeta sum has a pole at the incident point $u=u_{\ell_{\mathrm{inc}}}$. Moreover, the definitions of $A(u)$, $B(u)$ and hence of $P_{1,3}(u)$ and $R(u)$ involve sums only over $I_{\mathrm{scat}}$, so $R(u)$ is analytic at $u=u_{\ell_{\mathrm{inc}}}$ as well. Therefore $Q_{\mathrm{scat}}$ is analytic at $u=u_{\ell_{\mathrm{inc}}}$, and hence, by the physical lift $u=u(\zeta)$ and the injectivity of the uniformization (21), analytic at $\zeta=\zeta_{i}$. ∎
@@ 12 Radiation condition and Meixner edge condition in the spectral formulation
$S_{\eta}:=\{\zeta\in\mathbb{C}:|\Im\zeta|<\eta\},$ (68)
@@
$|Q(\zeta)|+|S(\zeta)|\leq C(1+|\zeta|)^{N},\qquad\zeta\in S_{\eta}\setminus\{\text{poles}\}.$ (69)
@@
$Q_{\mathrm{scat}}(u_{0})=0,\qquad\text{equivalently }Q_{\mathrm{scat}}(\zeta)\to 0\ \text{as }\Im\zeta\to+\infty,$ (70)
(and likewise for $S_{\mathrm{scat}}$). Then the Sommerfeld integrals (5)–(6) define classical solutions of the Helmholtz equations in their respective sectors, satisfy the Sommerfeld radiation condition as $r\to\infty$, and satisfy the Meixner finite-energy condition at the wedge apex $r\to 0$.
@@
For $r\to\infty$, steepest descent on the phase $\cos z$ along the deformed contour yields an outgoing leading term proportional to $e^{\mathrm{i}kr}/\sqrt{r}$, with remainder $o(r^{-1/2})$; see, for example, Bleistein–Handelsman [17, Ch. 6] or Wong [18, §2.4]. The outgoing far-field expansion implies the Sommerfeld radiation condition.
For $r\to 0$, one expands $e^{\mathrm{i}kr\cos z}=1+O(r)$ uniformly on $\gamma$ and uses (70) together with strip analyticity to shift the contour upward, obtaining boundedness of $u$ and its first derivatives in a neighborhood of the apex; boundedness of $u$ and $\nabla u$ implies the Meixner finite-energy condition (see [6, §3] and [4, Ch. 3]). ∎
Corollary 12.2.
@@ 13 Far-field diffraction coefficient
The contour deformation and stationary-phase evaluation used in this section are justified under standard analyticity and growth hypotheses for $Q_{\mathrm{scat}}(\zeta)$ in a strip containing the real axis; see, for example, [17, §2.4] or [18, Ch. II]. In the present lemniscatic case these hypotheses are met for $\varepsilon>0$ because $Q_{\mathrm{scat}}$ is given by an explicit elliptic-function representation (Theorem 10.4) and the forcing poles are displaced off the real $\zeta$-axis by the limiting absorption parameter.
@@
Assume the analytic strip and growth framework of §1.1 and the uniqueness principle of Lemma 1.3. Then the diffracted far-field coefficient is
$D(\theta,\theta_{i})=e^{-\mathrm{i}3\pi/4}\sqrt{\frac{2}{\pi k_{0}}}\,Q_{\mathrm{scat}}(\theta),$ (71)
where $Q_{\mathrm{scat}}(\theta)$ denotes the physical boundary value of $Q_{\mathrm{scat}}(\zeta)$ at $\zeta=\theta$, obtained by evaluating the physical lift $u(\zeta)$ and taking the limiting absorption limit $\varepsilon\to 0^{+}$ at the end.
@@
$u^{(0)}_{\mathrm{scat}}(r,\theta)=\frac{1}{2\pi\mathrm{i}}\int_{\gamma}e^{\mathrm{i}k_{0}r\cos z}\,\big(Q_{\mathrm{scat}}(\theta+z)-Q_{\mathrm{scat}}(\theta-z)\big)\,\,\mathrm{d}z.$
@@
$u^{(0)}_{\mathrm{scat}}(r,\theta)=\frac{1}{2\pi\mathrm{i}}\left(\int_{\gamma}+\int_{-\gamma}\right)e^{\mathrm{i}k_{0}r\cos z}\,Q_{\mathrm{scat}}(\theta+z)\,\,\mathrm{d}z.$
Under the analyticity and strip-growth hypotheses stated above, the union $\gamma\cup(-\gamma)$ may be deformed to the steepest descent rays through the saddle $z=0$, where $\cos z=1-\frac{1}{2}z^{2}+O(z^{4})$. The leading contribution is therefore
$u^{(0)}_{\mathrm{scat}}(r,\theta)\sim\frac{2}{2\pi\mathrm{i}}\,Q_{\mathrm{scat}}(\theta)\,e^{\mathrm{i}k_{0}r}\int_{-\infty}^{\infty}\exp\!\left(-\mathrm{i}\frac{k_{0}r}{2}x^{2}\right)\,\mathrm{d}x,\qquad r\to\infty.$
Using $\int_{-\infty}^{\infty}e^{-\mathrm{i}ax^{2}}\,\mathrm{d}x=\sqrt{\pi/a}\,e^{-\mathrm{i}\pi/4}$ for $a>0$ yields
$u^{(0)}_{\mathrm{scat}}(r,\theta)\sim\frac{e^{\mathrm{i}k_{0}r}}{\sqrt{r}}\,D(\theta,\theta_{i}),\qquad D(\theta,\theta_{i})=e^{-\mathrm{i}3\pi/4}\sqrt{\frac{2}{\pi k_{0}}}\,Q_{\mathrm{scat}}(\theta),$
which is (71). This normalization is consistent with the standard two-dimensional GTD convention for wedge diffraction [3]. ∎
@@ 14 Conclusion and outlook
  • A Sommerfeld spectral representation for the transmission problem is reduced to a closed two-face functional system in the spectral variable (§1.1).
@@
  • The jet-killing polynomials $p,q$ are constructed so that $A(u)+p(t(u))=O(t(u)^{4})$ and $B(u)+q(t(u))=O(t(u)^{4})$ on $\Omega_{\mathrm{phys}}^{+}$, yielding analyticity at the physical basepoint $u_{0}$ (Section 8).
@@
  • Reciprocity of the extracted far-field coefficient is not established; numerical tests indicate a non-reciprocal coefficient for generic angles, so the physical validity of the closed form is not claimed.
@@ 15 Symbolic evaluation recipe
    Define $u_{\ell}$ (physical lift) by the uniformization (32).
@@
    Compute $\alpha_{\ell},\beta_{\ell},C_{\ell}$ from Propositions 6.16.3.
@@
    Compute jet coefficients $(p_{1},p_{2},p_{3})$ and $(q_{1},q_{2},q_{3})$ from (49)–(50).
@@
    Build $A(u)$ and $B(u)$ from (47).
@@
    Build $P_{1,3}(u)$ and compute $d_{\ell}$ from (65) or Proposition 10.1. Then form $R(u)$ via (66).
@@
    For a given $\zeta$, compute $(t(\zeta),Y(\zeta))$ from the physical branch of $s(t,Y)=s_{\zeta}$ (19), lift to $u(\zeta)$ via (21), and evaluate $Q_{\mathrm{scat}}(\zeta)=Q_{\mathrm{scat}}(u(\zeta))$.
@@
    Obtain the far-field diffraction coefficient from (71).